An engine supplies 196 joules of energy. If the energy is supplied to a weight of 500 gms. How


high can it be lifted


(a) 38.2


(b) 39.2


(c) 40.2


(d) 42

Grade:11

2 Answers

Hrishant Goswami
10 years ago

Ans- (b) Energy supplied to the engine = 196 JMass of water = 500 gm = 500/1000 = ½ kg2Acceleration due to ground (g) = 10 mt/sec .Energy required for lifting water = mghH = energy supplied / m x g = 196 x 2 / 1 x 10 = 39.2 mt.

JUnnarkar
10 years ago

39.2

Energy supplied to the engine = 196 J 

Mass of water = 500 gm = 500/1000 = ½ kg

 Acceleration due to ground (g) = 10 mt/sec2. 

Energy required for lifting water = mgh H = energy supplied / m x g = 196 x 2 / 1 x 10 = 39.2 mt.